OpenStudy (anonymous):

How many kg of pure salt must be added to 8 kg of a 30% solution to obtain a 35% soltution? Round to the nearest tenth.

OpenStudy (amistre64):

8(.30-.35)/(1-.35); maybe, and drop the sign

OpenStudy (anonymous):

Thanks OOOOOH

OpenStudy (amistre64):

is it right? cause I cant tell yet

OpenStudy (anonymous):

I don't know

OpenStudy (amistre64):

well, its help if you knew how to know :) let me dbl chk it

OpenStudy (amistre64):

a(-.35) + 8(-.35) = b(-.35) a(1) + 8(.30) = b(.35) ------------------------- a(1-.35) + 8(.30-.35) = 0 a = -8(.30-.35)/(1-.35)

OpenStudy (amistre64):

it should be good, first one goes negative but i knew that lol

OpenStudy (amistre64):

8/13 i believe is what the simplifies to